Are there values of (a) belongs to C for which the matrix [
1 0 0
0 -1/√2 1/√2
0 1/√2. a
]
is unitary? Justify your answer
1
Expert's answer
2020-12-01T15:52:46-0500

"\displaystyle\textbf{\textsf{Yes.}}\\

\textsf{Let}\, A = \begin{pmatrix}
1 & 0 & 0\\
0 & -\frac{1}{\sqrt{2}} & \frac{1}{\sqrt{2}} \\
0 & \frac{1}{\sqrt{2}} & a
\end{pmatrix} = \begin{pmatrix}
1 + 0i & 0 + 0i & 0 + 0i\\
0 + 0i & -\frac{1}{\sqrt{2}} + 0i & \frac{1}{\sqrt{2}} + 0i\\
0 + 0i & \frac{1}{\sqrt{2}} + 0i & a + 0i
\end{pmatrix}\\


\textsf{Conjugate of}\, A = A^{c} = \begin{pmatrix}
1 & 0 & 0\\
0 & -\frac{1}{\sqrt{2}} & \frac{1}{\sqrt{2}} \\
0 & \frac{1}{\sqrt{2}} & a
\end{pmatrix} =\begin{pmatrix}
1 - 0i & 0 - 0i & 0 - 0i\\
0 - 0i & -\frac{1}{\sqrt{2}} - 0i & \frac{1}{\sqrt{2}} - 0i\\
0 - 0i & \frac{1}{\sqrt{2}} - 0i & a - 0i
\end{pmatrix} = \begin{pmatrix}
1 & 0 & 0\\
0 & -\frac{1}{\sqrt{2}} & \frac{1}{\sqrt{2}} \\
0 & \frac{1}{\sqrt{2}} & a
\end{pmatrix}\\


\textsf{Conjugate transpose of}\, A = A^{\ast} = \begin{pmatrix}
1 & 0 & 0\\
0 & -\frac{1}{\sqrt{2}} & \frac{1}{\sqrt{2}}\\
0 & \frac{1}{\sqrt{2}}& a
\end{pmatrix}\\



AA^{\ast} = A^2 = \begin{pmatrix}
1 & 0 & 0\\
0 & 1 & \frac{a}{\sqrt{2}} - \frac{1}{2}\\
0 & \frac{a}{\sqrt{2}} - \frac{1}{2}& a^2 + \frac{1}{2}
\end{pmatrix}\\


\textsf{By definition}\, AA^{\ast} = \begin{pmatrix}
1 & 0 & 0\\
0 & 1 & 0\\
0 & 0 & 1
\end{pmatrix}\\


\textsf{Comparing both elements}\\


a^2 + \frac{1}{2} = 1 \\

a^2 = \frac{1}{2} \\

a = \pm\frac{1}{\sqrt{2}} = \pm\frac{\sqrt{2}}{2}\\


\frac{a}{\sqrt{2}} - \frac{1}{2} = 0\\

\frac{a}{\sqrt{2}} = \frac{1}{2} \\

a = \frac{\sqrt{2}}{2}\\


\textsf{The positive value of}\, a \,\,\textsf{satisfies}\\
\textsf{the definition of a unitary matrix}.\\
\textsf{The positive value satisfies the}\\

\textsf{identity matrix because on}\\

\textsf{substituting the negative value of}\, a, \\

\textsf{we do not have an identity matrix,}\\

\textsf{the}\, 1's \, \textsf{on substitution of}\\

\textsf{the positive value}\,a\, \textsf{becomes}\, -1\\

\textsf{on the substitution of the negative value}\\

\textsf{of}\,a.\\

\textsf{Also,}\,\,\frac{\sqrt{2}}{2}\in \mathbb{C}."